OpenStudy (anonymous):

solve this by the addition method 5x+4y=12 7x-6y=40

OpenStudy (saifoo.khan):

x = 4 y = -2

OpenStudy (anonymous):

Is that the addition method?:-)

OpenStudy (anonymous):

Just apply what I taught you in the last question, lol. Try to multiply the first equation and the second equation by two different values so that when added they result in 0x.

OpenStudy (anonymous):

Think you can manage that, Eva147?

OpenStudy (anonymous):

Good for u.

OpenStudy (anonymous):

Maybe you need another example .... Multiply the first equation through by 7 and the second equation through by -5 35x+28y=84 -35x+30y=-200 ++++++++++++ Add 0x+58y=-116 y=-2 Substitute that into either equation to get the value of x and check with the other equation to make sure it works out.